OpenStudy (jdanmath):

Binomial distribution: n=19, p=0.07, q= 0.93 x=2 or more how do I find the probability of x being 2 or more?

OpenStudy (holsteremission):

Given that \(X\sim\mathcal B(n,p)\), you have \[\mathbb P(X\ge2)=\sum_{i=2}^n\mathbb P(X=i)=\sum_{i=2}^n\binom nip^{n-i}q^i\]You can do less work by computing the complement probability: \[\begin{align*} \mathbb P(X\ge k)&=1-\mathbb P(X<k)\\[1ex] \implies \mathbb P(X\ge2)&=1-\mathbb P(X<2)\\[1ex] &=\mathbb P(X=0)+\mathbb P(X=1) \end{align*}\]where \[\mathbb P(X=k)=\binom nkp^{n-k}q^k\]